ilog (x-i/x+i) should be equal to what value pls give me the answer

ilog  (x-i/x+i) should be equal to what value pls give me the answer
 

Grade:12

1 Answers

Sumit
5 years ago
blog(x-i/x+i.x-i/x-i)
I log{(x-i)2/x2-i2)}
ilog{(x-i)2/x2+1)}.           (i2=-1)
i{log(x-i)2-log(x2+1)}.      (loga/b=loga-logb)
i{2log(x-i)-log(x2+1)}.       ( logab=bloga)
2ilog(x-i)-ilog(x2+1).  Answer👍